[The following information applies to the questions displayed below.]

Adger Corporation is a service company that measures its output based on the number of customers served. The company provided the following fixed and variable cost estimates that it uses for budgeting purposes and the actual results for May as shown below:

Fixed Element
per Month
Variable Element per Customer Served Actual Total
for May
Revenue $ 6,200 $ 196,000
Employee salaries and wages $ 69,000 $ 1,600 $ 122,900
Travel expenses $ 610 $ 17,700
Other expenses $ 48,000 $ 45,200

When preparing its planning budget the company estimated that it would serve 30 customers per month; however, during May the company actually served 35 customers.

1. What amount of revenue would be included in Adger’s flexible budget for May?

2. What amount of employee salaries and wages would be included in Adger’s flexible budget for May?

3. What amount of travel expenses would be included in Adger’s flexible budget for May?

4. What amount of other expenses would be included in Adger’s flexible budget for May?
Answer #1

1) Revenue in flexible budget = 6200*35 = 217000

2) Employee salaries and wages = (1600*35+69000) = 125000

3) Travel expense = 610*35 = 21350

4) Other expense = 48000